# Patch2Loc: Learning to Localize Patches for Unsupervised Brain Lesion Detection

### v1

@@ Abstract
Detecting brain lesions as abnormalities observed in magnetic resonance imaging (MRI) is essential for diagnosis and treatment. In the search of abnormalities, such as tumors and malformations, radiologists may benefit from computer-aided diagnostics that use computer vision systems trained with machine learning to segment normal tissue from abnormal brain tissue. While supervised learning methods require annotated lesions, we propose a new unsupervised approach (Patch2Loc) that learns from normal patches taken from structural MRI. We train a neural network model to map a patch back to its spatial location within a slice of the brain volume. During inference, abnormal patches are detected by the relatively higher error and/or variance of the location prediction. This generates a heatmap that can be integrated into pixel-wise methods to achieve finer-grained segmentation. We demonstrate the ability of our model to segment abnormal brain tissues by applying our approach to the detection of tumor tissues in MRI on T2-weighted images from BraTS2021 and MSLUB datasets and T1-weighted images from ATLAS and WMH datasets. We show that it outperforms the state-of-the art in unsupervised segmentation. The codebase for this work can be found on our GitHub page.
Index Terms:
Anomaly detection, Convolutional neural networks, Computer aided diagnosis, Machine learning, Magnetic resonance imaging, Neuropathology, Neuroradiology, Self-supervised learning, Tumors
Figure 1: Schematic diagram for Patch2Loc. (Top) Training stage: A randomly selected patch from a normal MRI slice (patches with less than $20\%$ of its content as brain tissues, as in the yellow patch, are rejected) is paired with its two-dimensional location $Y$ and slice index $A$. The relationship between the image patch, slice, and its location is modeled as a conditional distribution $P_{\theta}(Y|X,A)$ using a 2D Gaussian distribution defined by the mean the variance of each coordinate that are functions of the patch and slice index. (Bottom) Inference stage: Overlapping patches with a fixed stride are extracted as in convolution from an MRI slice and the model is applied to each patch. The norm of the difference between the model’s predicted mean and the patch’s true location creates an error map. Likewise, the sum of the variances create a variance map. Together the norm of the error and variance can identify anomalous patches.
I Introduction
Detecting and localizing abnormal brain tissue through non-invasive neuroimaging is a critical medical task with profound implications for patients’ well-being and overall quality of life. Early detection and intervention can mitigate severe consequences in cases such as incipient brain tumors or cortical malformations, which are associated with drug-resistant pediatric epilepsy [1]. The intricate structural complexity of the brain necessitates the expertise of neurologists and radiologists for accurate diagnosis. The development of computer-aided diagnostics at an expert level is motivated by the shortage of neuroradiologists [2].
Magnetic resonance imaging (MRI) has emerged as a reliable modality for detecting lesions within the brain [3, 4, 5]. MRI offers the distinct advantage of being radiation-free and non-invasive and different sequences can produce various tissue contrasts. In recent years, significant strides have been made in leveraging machine learning, particularly deep learning techniques for computer vision [6] to automate the detection and localization of abnormal brain tissues in neuroimages. Methods generate abnormality scores at the pixel or voxel-level that can be thresholded to segment images abnormal and normal regions. This progress is largely attributable to the availability of extensive datasets and advanced deep learning methodologies.
While models trained with supervision may be expected to exhibit superior performance, they require extensive annotations delineating the anomalies. However, the release of detailed annotations is limited, especially for rare abnormalities [7, 8]. This motivates the improvement of unsupervised learning methods that use only putative normal images that can then be applied for abnormal brain tissue detection. Possibly using a minimal set of labeled images for hyper-parameter selection.
As discussed in Section V, state-of-the-art methods use autoencoders, denoising autoencoders, and diffusion models learn to reconstruct or denoise normal images based on the typical structure, such that the reconstructions of abnormal portions of images is closer to typical structure and the reconstruction errors highlight abnormalities. In particular, denoising diffusion probabilistic models (DDPMs) have recently outperformed other generative and restoration-based approaches for unsupervised abnormality detection [9, 10, 11, 12, 13]. The key idea is to train a DDPM on normal brain MRI volumes (3D images) or slices (2D image taken from the 3D image of the brain). Since the model learns the patterns of normal brains, if it is fed with slices contain abnormal regions, it inpaints abnormal regions with normal patterns. The reconstruction error, typically measured as the pixel-wise difference between the original and reconstructed images, serves as a heatmap for potential anomalies, with higher errors indicating regions more likely to contain abnormalities.
Despite their effectiveness, DDPMs suffer from the noise paradox [9, 14], where the level of added noise during inference significantly impacts performance: higher noise levels remove more anomalies but increase false positives due to over-inpainting of normal tissues. This trade-off complicates the detection process and limits DDPM’s reliability in clinical settings. Furthermore, DDPMs and the vast majority of existing unsupervised abnormality segmentation methods for brain MRI rely on processing entire slices. Thus, these methods detect localized lesions from a “global” context. In contrast, analyzing local patterns in isolation, given their location, may provide more focused detection of abnormalities.
With this motivation, we introduce Patch2Loc, a novel framework tailored for achieving unsupervised abnormality segmentation in neuroimaging. We create a self-supervised task, predicting the location of a patch, such that given the true location the differences and uncertainty with the prediction can identify local anomalies.
I-A Main Contributions
Our approach offers several key advantages:
@@ I-A Main Contributions
    Spatially-Aware Local Feature Learning: Patch2Loc learns to predict the spatial origin of a given patch within the brain. By encoding local anatomical and tissue structure patterns from normal tissue patches, the model can effectively distinguish abnormal patches based on deviations in location prediction.
@@
    Uncertainty-Guided Anomaly Detection: The variance in Patch2Loc’s predictions serves as an uncertainty estimate, enabling more robust abnormality detection. By combining location error and uncertainty, we derive an enhanced abnormality score that improves detection performance.
@@
    Minimal Hyperparameter Dependence: Unlike DDPM-based methods, which require hyperparameter tuning (e.g., noise level selection), Patch2Loc is self-supervised and adaptable. The most anomalous patches can be found by the abnormality score.
By shifting the focus to local features and leveraging self-supervised learning, Patch2Loc’s abnormality score is intuitive and easy to interpret and avoidance of extensive hyperparameter tuning and labeled data increases clinical applicability.
II Methodology
The idea underlying Patch2Loc is that there is a strong relationship between patch content and location, due to the local patterns in structural neuroimages, that is used by neuroradiologists when identifying structural abnormalities. With sufficient sampling across the population, this relationship can be modeled using machine learning. As each brain has slightly different sizes, the images are first registered using rigid transformation. Fig. 1 shows a summary schematic diagram for methodology behind Patch2Loc.
II-A Problem Formulation
Let $(Y_{1},Y_{2},A)\in[0,100]^{3}$ denote the 3D location of a rectangular image patch $X\in\mathcal{X}\subset\mathbb{R}_{\geq 0}^{S_{1}\times S_{2}}$ within the brain taken at the 2D location $Y=(Y_{1},Y_{2})\in[0,100]^{2}$ from the slice located at $A\in[0,100]$. The coordinates are percentages of the patch’s location in absolute coordinates $(L_{1},L_{2},L_{3})$ relative to the brain volume’s spatial extent along each axis $(E_{1},E_{2},E_{3})$: $Y_{1}=100\cdot\frac{L_{1}}{E_{1}}$, ${Y}_{2}=100\cdot\frac{L_{2}}{E_{2}}$, and ${A}=100\cdot\frac{L_{3}}{E_{3}}$. With registered brain scans, the spatial extents $E_{1},E_{2},E_{3}$ are constant for all scans. Each 2D patch is rectangular with an absolute size of $S_{1}\times S_{2}$ chosen based on a fixed relative proportion $r=\frac{S_{1}}{E_{1}}=\frac{S_{2}}{E_{2}}$. The choice of $r$ controls the patch’s coverage.
@@ II-A Problem Formulation
The functions describing the conditional mean and variances $\mu^{\theta}$ and $\Sigma^{\theta}$ are modeled using a neural network with parameters $\theta=(\theta_{0},\theta_{m},\theta_{v})$, where $\theta_{0}$ are the parameters of the shared patch encoder $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$, $\theta_{m}$ are the parameters of the mean prediction head $\tilde{\mu}$, such that $\mu^{\theta}(X,A)=\tilde{\mu}(\phi(X),A)\in\mathbb{R}^{2}$, and $\theta_{v}$ are the parameters particular to the log-variance head $\tilde{\varsigma}$, such that $\varsigma^{\theta}(X,A)=\tilde{\varsigma}(\phi(X),A)\in\mathbb{R}^{2}$ and $\Sigma_{ii}^{\theta}(X,A)=\mathrm{exp}(\varsigma_{i}^{\theta}(X,A)),\quad i\in\{1,2\}$.
II-B Loss Function
@@
$-\log P_{\theta}(Y|X,A)\propto\sum_{i=1}^{2}\frac{\lvert Y_{i}-\mu_{i}^{\theta}(X,A)\rvert^{2}}{\mathrm{exp}(\varsigma_{i}^{\theta}(X,A))}+\varsigma_{i}^{\theta}(X,A),$ (3)
which causes convergence to local minima with poor performance [15, 16]. Hence, we adapt the $\beta$-NLL loss proposed in the work by [16] that scales each dimension of the loss by the variance estimate taken to the $\beta$th power, $\beta\in(0,1]$,
@@
[…]
II-C Abnormality Scores
@@
To detect abnormalities with Patch2Loc, we calculate an abnormality score from both location prediction error and the variance estimates, after log transformation. The log of the squared error is calculated as $\log(\lVert Y-\mu^{\theta}(X,A)\rVert_{2})^{2}$. The log-variance is calculated as $\log(\mathrm{exp}(\varsigma_{1}^{\theta}(X,A))+\mathrm{exp}(\varsigma_{2}^{\theta}(X,A)))$. We define the abnormality score as the sum of the log-squared error and log-variance, which is the $\beta$-NLL loss (4) for $\beta=1$.
II-D Model Details
@@
While the Patch2Loc task can be applied to different anatomical orientations for the slices, we consider axial slices, such that $A$ is proportional to the location of the slice along the vertical dimension (bottom to top of the brain). We select $r=12.5\%$ for the proportion of the axial slice’s width and length. Ideally, the smaller the patch, the more precise the resulting heatmap would be. However, in development we found that smaller patch sizes deteriorates Patch2Loc’s performance since the patches are more ambiguous. When the patch size is too small, the inherent structural symmetry of the brain causes patches from different locations to appear similar, making it challenging for Patch2Loc to distinguish between them effectively.
@@
For the encoding model $\phi:\mathbb{R}_{\geq 0}^{S_{1}\times S_{2}}\rightarrow\mathbb{R}^{d}$, which serves as a common backbone, we use ResNet-18 [17], producing a shared latent feature vector with dimension $d=512$. Since the patch size are small (i.e., $S_{1}\times S_{2}=24\times 24$), we replace the ResNet-18 first convolution layer with a kernel of size 3 instead of 7. The latent vector from the backbone $\phi(X)\in\mathbb{R}^{d}$ is added to a sinusoidal positional encoding of the slice coordinate $A$ [18]. This representation is input to two separate branches for the mean $\tilde{\mu}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2}$ and the log-variance $\tilde{\varsigma}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2}$. Each branch consists of four fully-connected (FC) layers with output dimensions $128,64,32$, and a batch normalization layer and a rectified linear unit (ReLU) activation function between each two linear layers, followed by a linear layer to the two-dimensional outputs. For the $\beta$-NLL loss in (4), we choose the suggested value of $\beta=0.5$ [16]. We confirmed that using NLL as a loss directly results in a local minimum where the model produces a high variance estimate and the mean has a high error.
II-E Convolutional Approach for Pixel-level Detection
@@
To perform pixel-level abnormality detection, we employ Patch2Loc in a convolutional manner. Specifically, for each slice we feed overlapping patches with a specified stride into the Patch2Loc model. Subsequently, we utilize the Patch2Loc outputs, predicted mean and variance, along with the maps of the true coordinates sampled at the same stride to construct a scoring map. A stride of 1 results in a scoring map identical in size to the testing MRI image, albeit with increased computational demands. This is depicted in the bottom section of Fig. 1.
[…]
III Data

### v2

@@ Abstract
Detecting brain lesions as abnormalities observed in magnetic resonance imaging (MRI) is essential for diagnosis and treatment. In the search of abnormalities, such as tumors and malformations, radiologists may benefit from computer-aided diagnostics that use computer vision systems trained with machine learning to segment normal tissue from abnormal brain tissue. While supervised learning methods require annotated lesions, we propose a new unsupervised approach (Patch2Loc) that learns from normal patches taken from structural MRI. We train a neural network model to map a patch back to its spatial location within a slice of the brain volume. During inference, abnormal patches are detected by the anomaly score based on the error and variance of the location prediction. By applying the network in a convolutional manner, this generates a pixel-wise heatmap of anomalies providing finer-grained segmentation. We demonstrate the ability of our model to segment abnormal brain tissues by applying our approach to the detection of tumor tissues in MRI on T2-weighted images from BraTS2021 and MSLUB datasets and T1-weighted images from ATLAS and WMH datasets. We show that it outperforms the state-of-the art in unsupervised segmentation.
1 INTRODUCTION
Detecting and localizing abnormal brain tissue in neuroimages is a critical diagnostic task, where early intervention can mitigate severe outcomes like those from incipient tumors or cortical malformations associated with epilepsy (Josephson et al., 2012). While magnetic resonance imaging (MRI) is a powerful non-invasive modality for this task, modern supervised deep learning models are hindered by the scarcity of annotated data, particularly for rare and structurally diverse conditions (Busby et al., 2018; Hagens et al., 2019). This data bottleneck, combined with a shortage of neuroradiology experts (Merewitz and Sunshine, 2006), makes the development of robust unsupervised methods for anomaly detection an essential clinical and research goal.
Current state-of-the-art unsupervised approaches, including autoencoders (Zimmerer et al., 2018; Sato et al., 2019; Meissen et al., 2022b, a; Behrendt et al., 2022; Bercea et al., 2025), adversarial autoencoders (Brock et al., 2018; Chen and Konukoglu, 2018; Baur et al., 2019), denoising autoencoders (Kascenas et al., 2022), transformers (Pinaya et al., 2022a) and denoising diffusion probabilistic models (DDPMs) (Wyatt et al., 2022; Pinaya et al., 2022b; Liang et al., 2023; Bercea et al., 2025; Behrendt et al., 2024, 2025), operate on a principle of global reconstruction. These methods learn the typical structure of a normal brain and flag anomalies as regions where the model fails to reconstruct the input accurately. However, this reliance on global context might provide enough clues to reconstruct the abnormal regions (Wyatt et al., 2022; Bercea et al., 2023c; Kascenas et al., 2023). The best-performing models are DDPMs, but they suffer from a difficult trade-off called the ‘noise paradox’ (Kascenas et al., 2023; Bercea et al., 2023a)—where enhancing anomaly signals can degrade the reconstruction of normal tissue, leading to false positives and limiting clinical reliability. These methods are thus constrained by a delicate balancing act of an inference-time hyperparameter (i.e., the amount of added noise) to be effective.
To overcome these limitations, we propose Patch2Loc, a novel framework that fundamentally shifts the paradigm from global reconstruction to local structural assessment. Instead of learning to reconstruct an entire image, Patch2Loc is trained on a simple yet powerful localizing task: predicting the spatial coordinates of an isolated image patch from a normal brain. The intuition is that the model learns the distinct anatomical patterns characteristic of each location. When presented with a patch containing an anomaly, its structure deviates from the norm, causing the model to predict its location with high error and high uncertainty. This provides a direct and robust signal for anomaly detection based on localized structural failure. This is a key distinction from other location-based self-supervised learning (SSL) tasks such as Jigsaw (Noroozi and Favaro, 2016), patch ordering and context prediction (Doersch et al., 2015; Taleb et al., 2020), where the objective is to learn from the relationships between multiple patches (e.g., context or relative position). These tasks use the relationship between two or more patches within the global context. Their task losses are not localized to a single patch; thus, they were not designed for localized abnormality detection. In contrast, Patch2Loc can be applied convolutionally using overlapping patches to localize abnormalities by increases in the location prediction error and the model’s uncertainty.
In summary, Patch2Loc offers several key advantages:
@@ 1 INTRODUCTION
    Spatially-Aware Local Feature Learning: By learning to predict a patch’s spatial origin, Patch2Loc encodes local anatomical patterns. This allows it to effectively distinguish abnormal patches based on deviations in location prediction, directly targeting the source of the anomaly.
@@
    Uncertainty-Guided Anomaly Detection: The variance in the model’s predictions serves as a powerful uncertainty estimate. Combining the location prediction error with this uncertainty creates a abnormality score that better correlates with ground truth annotations.
@@
    Minimal Inference Hyperparameter Dependence: Unlike reconstruction-based methods that require careful tuning of noise levels during inference, Patch2Loc operates with a stable configuration, providing a pixel-level abnormality heatmap without complex post-training adjustments.
By focusing on local features and leveraging both prediction error and uncertainty, Patch2Loc offers an intuitive, interpretable, and effective solution for unsupervised anomaly detection with high potential for clinical applicability.
Figure 1: Schematic diagram for Patch2Loc. (Top) Training stage: A randomly selected patch from a normal MRI slice is paired with its two-dimensional location $Y$ and slice index AA. The relationship between the image patch, slice, and its location is modeled as a conditional distribution $P_{\theta}(Y|X,A)$ using a 2D Gaussian distribution defined by the mean and the variance of each coordinate that are functions of the patch and slice index. Note that patches with less than $20\%$ of its content as brain tissues, as in the yellow patch, are rejected. (Bottom) Inference stage: Overlapping patches with a fixed stride are extracted as in convolution from an MRI slice and the model is applied to each patch. The squared norm of the error between the model’s predicted mean and the patch’s true location creates an error map. Likewise, the sum of the variances create a variance map. Together the sum of the logarithms of the errors and variances highlight anomalous areas.
2 METHODOLOGY
The idea underlying Patch2Loc is that there is a strong relationship between patch content and location, due to the normal anatomical patterns in structural neuroimages, that is used by neuroradiologists when identifying structural abnormalities. With sufficient sampling across the population, this relationship can be modeled using machine learning. As each brain has slightly different sizes, the images are first registered using rigid transformation, such that the patches taken from the same coordinates contain similar anatomy. Figure 1 shows a summary schematic diagram for methodology behind Patch2Loc.
2.1 Problem Formulation
Let $(Y_{1},Y_{2},A)\in[0,100]\times[0,100]\times[0,100]$ denote the 3D location of a rectangular image patch $X\in\mathcal{X}\subset\mathbb{R}_{\geq 0}^{S_{1}\times S_{2}}$ within the brain taken at the 2D location $Y=(Y_{1},Y_{2})\in[0,100]\times[0,100]$ from the slice located at $A\in[0,100]$. The coordinates are percentages of the patch’s location in absolute coordinates $(L_{1},L_{2},L_{3})$ relative to the brain volume’s spatial extent along each axis $(E_{1},E_{2},E_{3})$: $Y_{1}=100\cdot\frac{L_{1}}{E_{1}}$, ${Y}_{2}=100\cdot\frac{L_{2}}{E_{2}}$, and ${A}=100\cdot\frac{L_{3}}{E_{3}}$. With registered brain scans, the spatial extents $E_{1},E_{2},E_{3}$ are constants common to all scans. Each 2D patch is rectangular with an absolute size of $S_{1}\times S_{2}$ chosen based on a fixed relative proportion $r=\frac{S_{1}}{E_{1}}=\frac{S_{2}}{E_{2}}$. The choice of $r$ controls the patch’s coverage.
@@ 2.1 Problem Formulation
The functions predicting the conditional mean and variances $\mu^{\theta}$ and $\Sigma^{\theta}$ of location given the patch are modeled using a neural network with parameters $\theta=(\theta_{0},\theta_{m},\theta_{v})$, where $\theta_{0}$ denotes the parameters of the shared patch encoder $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$, $\theta_{m}$ denotes the parameters of the mean prediction head $\tilde{\mu}$, such that $\mu^{\theta}(X,A)=\tilde{\mu}(\phi(X),A)\in\mathbb{R}^{2}$, and $\theta_{v}$ denotes the parameters of log-variance head $\tilde{\varsigma}$, such that $\varsigma^{\theta}(X,A)=\tilde{\varsigma}(\phi(X),A)\in\mathbb{R}^{2}$ and $\Sigma_{ii}^{\theta}(X,A)=\mathrm{exp}(\varsigma_{i}^{\theta}(X,A)),\quad i\in\{1,2\}$.
2.2 Loss Function
@@
$\log P_{\theta}(Y|X,A)\propto\sum_{i=1}^{2}\frac{\lvert Y_{i}-\mu_{i}^{\theta}(X,A)\rvert^{2}}{\mathrm{exp}(\varsigma_{i}^{\theta}(X,A))}+\varsigma_{i}^{\theta}(X,A),$ (3)
[…]
@@
where $\lfloor\cdot\rfloor$ indicates the stop-gradient operation such that gradients are not taken with respect to this scaling. This scaling mitigates the effect of the variance estimate on the gradient of the loss with respect to the mean. Notably, when $\beta=1$ then $\mu_{i}^{\theta}$ is updated based on mean-squared error. Thus, we choose the suggested value of $\beta=0.5$ (Seitzer et al., 2022).
2.3 Abnormality Score
@@
To detect abnormalities with Patch2Loc, we calculate an abnormality score from the location prediction error and the sum of the variance estimates, after a log transformation. The log-error is computed as $\log(\lVert Y-\mu^{\theta}(X,A)\rVert_{2}^{2}+\varepsilon)$, where $\varepsilon=0.5$ such that it does not effect large errors. The log-variance is computed as $\frac{1}{2}\log(\mathrm{det}(\Sigma^{\theta}(X,A)))=\frac{1}{2}\sum_{i=1}^{2}\log(\Sigma^{\theta}_{ii}(X,A)))=\frac{1}{2}\sum_{i=1}^{2}\varsigma_{i}^{\theta}(X,A)$. The log-variance captures the model’s uncertainty since adding $1+\log(2\pi)$ to the log-variance is the entropy of $\mathcal{N}(\mu^{\theta}(X,A),\Sigma^{\theta}(X,A))$. The abnormality score is
$\begin{split}\textrm{Score}(X,Y,A)&=\underbrace{\log(\lVert Y-\mu^{\theta}(X,A)\rVert_{2}^{2}+\varepsilon)}_{\text{Log(Error\textsuperscript{2})}}\\ &\quad+\underbrace{\frac{1}{2}\varsigma_{1}^{\theta}(X,A)+\frac{1}{2}\varsigma_{2}^{\theta}(X,A)}_{\text{Log(Variance)}}.\end{split}$ (5)
It assigns the highest values to data points that are both far from the predicted mean and are associated with a large predicted variance. Essentially, when the model expects being wrong and is. Normal patches where the model is certain and have low error predictions have the lowest anomaly scores.
2.4 Model Details
@@
While the Patch2Loc task can be applied to different anatomical orientations for the slices, we consider axial slices, such that $A$ is indicates the location of the slice along the vertical dimension (bottom to top of the brain). We select $r=12.5\%$ for the proportion of the axial slice’s width and length. Ideally, the smaller the patch, the more precise the resulting heatmap would be. However, in development we found that smaller patch sizes deteriorates Patch2Loc’s performance since the patches are more ambiguous. That is, when the patch size is too small, the inherent structural symmetry of the brain causes patches from different locations to appear similar, making it challenging for Patch2Loc to distinguish between them effectively. Contrastingly, larger patches may fail to be useful for identifying anomalies because the model can use the context surrounding an abnormality to accurately predict a location.
@@
For the encoding model $\phi:\mathbb{R}_{\geq 0}^{S_{1}\times S_{2}}\rightarrow\mathbb{R}^{d}$, which serves as a common backbone, we use ResNet-18 (He et al., 2016), producing a shared latent feature vector with dimension $d=512$. Since the patch size are small (i.e., $S_{1}\times S_{2}=24\times 24$), we replace the ResNet-18 first convolution layer with a kernel of size 3 instead of 7. The latent vector from the backbone $\phi(X)\in\mathbb{R}^{d}$ is added to a sinusoidal positional encoding of the slice coordinate $A$ (Vaswani et al., 2017). This representation is input to two separate branches for the mean $\tilde{\mu}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2}$ and the log-variance $\tilde{\varsigma}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2}$. Each branch consists of four fully-connected layers with output dimensions $512,128,64,32$, with a batch normalization layer and then a rectified linear unit (ReLU) activation function after each fully-connected layer, followed by a final linear layer to the two-dimensional outputs.
2.5 Convolutional Approach for Pixel-level Detection
@@
To perform pixel-level abnormality detection, we employ Patch2Loc in a convolutional manner. Specifically, we feed overlapping patches with a specified stride into the Patch2Loc model, padding patches with zeros when necessary. Subsequently, we utilize the Patch2Loc outputs, predicted mean and variance, along with the corresponding true coordinates to construct the error, variance, and abnormality score map. This is depicted in the bottom section of Figure 1. When the stride is 1, the scoring map has the same resolution as the input MRI image. If a larger stride is used, the abnormality score heatmap will have a lower resolution, but it can be upsampled (e.g., via interpolation) to match the input dimensions. Throughout all experiments, we use a stride of 1.
3 RELATED WORK
While not used for unsupervised abnormality segmentation, one prior work by Taleb et al. (2020) uses self-supervised learning with a task that resembles Patch2Loc’s location prediction task. The task is a 3D version of a previous context-dependent self-supervised task proposed by Doersch et al. (2015). Specifically, the task consists of predicting the discrete relative location of a 3D patch among the possible locations in a $3\times 3\times 3$ grid surrounding a center patch that is provided as input. The performance for unsupervised abnormality segmentation was not benchmarked as the task served only as a pretext task to pretrain a backbone network before subsequent supervised segmentation (Taleb et al., 2020). Patch2Loc is distinguished in that its training directly informs the abnormality score. Additionally, Patch2Loc predicts the continuous location of a patch without the need of a center patch as context, using only the slice index as sufficient context for registered brain images.
Unsupervised abnormality segmentation in neuroimaging has benefited from improving machine learning models. Nonetheless, a notable baseline is based on a simple threshold on the image histogram (Meissen et al., 2022a), which can exceed the performance of baseline methods for certain modalities. Here we describe works relevant to Patch2Loc or those we compare against, a more comprehensive discussion of the related work can be found in Appendix A.
Many approaches leverage AE or VAE in novel ways. Rather than relying on squared error for training, another work by Meissen et al. (2022b) uses an AE architecture that reconstructs features obtained from a pretrained encoder using the Structural Similarity Index Measure (SSIM) as the loss function. SVAE by Behrendt et al. (2022) uses a VAE with transformers to capture the inter-slice dependencies and showed it can improve the results compared to 2D vanilla VAE. The RA method by Bercea et al. (2023c) uses an VAE with a cyclic loss and use the reconstruction error as abnormality score. Baur et al. (2021) noted that AE and VAE methods suffer from blurry reconstructed images, which hinders their performance. Incorporating a discriminator as in a GAN can improve the reconstruction quality.
One prior work by Van Hespen et al. (2021) used a patch-based auto-encoder with a cycle consistency term and a discriminator to distinguish between the original image and its reconstruction, testing it on specific abnormal tissues. We identified it as the only approach leveraging local features for unsupervised abnormality segmentation in brain MRI. However, the method was only applied to a dataset of infarcts (dead brain tissue caused by loss of blood flow) and was not compared for other abnormal tissues.
Another approach is to iteratively restore an image to better match the normal data distribution. PHANEs (Bercea et al., 2023b) uses a model to restore part of an MRI slice flagged by the RA method (Bercea et al., 2023c) to mitigate false positives within the flagged region. This is along the lines of denoising autoencoder (DAE) (Kascenas et al., 2022), which learns to remove correlated noise added to the input images during training. DAE outperformed GAN and VAE approaches, achieving higher Dice score and average precision.
Denoising diffusion probabilistic models (DDPM) (Ho et al., 2020) build on DAEs. One of the first works to use DDPM for unsupervised abnormality segmentation also proposed learning to denoise simplex noise instead of Gaussian noise to enhance performance (Wyatt et al., 2022). While this empirically works, how this matches the fundamental assumptions of diffusion processes is not clear. Specifically, simplex noise is procedurally generated, and is not described by a random process. In contrast, a diffusion model’s forward process is described by a Markov chain, with a Markov transition kernel (Sohl-Dickstein et al., 2015), typically described by adding Gaussian noise. Nonetheless, the simplex noise version of DDPM formulation essentially defines a training regime for denoising across different signal-to-noise ratios, where longer times correspond to higher noise regimes. During inference, denoising is performed from the initially high noise regime, then new simplex noise is applied at decreasing noise levels, and the denoising process continues. The patched diffusion model (pDDPM) by Behrendt et al. (2024) adds noise only to a patch of the whole input slice. The rest of the slice gives context for the DDPM to attempt to denoise the noised patch. Different versions of each slice with patches at different locations are used to identify abnormalities within the slice. Behrendt et al. (2025) used conditional DDPM (cDDPM) to denoise a slice given the latent embedding from a masked auto-encoder (MAE) (He et al., 2022) pretrained on normal MRI slices, which is further fine-tuned during training.
4 DATA